OpenStudy (anonymous):

A and B have coordinates –2 and 10 on a number line. If it is marked off into 10 congruent segments, which one of the following coordinates is NOT one of the marked points? a. –0.8 c. 1.6 b. –0.6 d. 2.8

OpenStudy (anonymous):

10-(-2)=12, which is the "lenght" of that number line. since you divide it by 10, each will be 12/10=1.2 you can do it the long way by adding 1.2 to -2 continuously and find out, for example -2+1.2=0.8, so it can't be A, etc. Or you can do it the shorter way but need more calculation, for example [-0.6-(-2)]/1.2 = -2.6/1.2 = 13/6 (false) So, the answer is c :)

OpenStudy (anonymous):

-2+1.2=-0.8, sorry I missed the " - " there

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

you are welcome :)